OpenStudy (anonymous):

Find the inverse of the following function 3/24x-13 Again, I know you have to swap X and Y values, but I ended up getting... 8x+ 104= y Something went wrong XD Please help

OpenStudy (anonymous):

I swear you need help with ever math problem there is XD but umm.. lemme try .-.

OpenStudy (anonymous):

3/24x-13=y Swap x and y 3/24y-13=x solve for y

OpenStudy (anonymous):

I know how to do inverse. I think I did something wrong, So I'm asking fo rhelp, I don't see what is wrong with clarifying my mistakes.

OpenStudy (anonymous):

That's what I tried to do tim and the answer I got was horrendous I think I messed something up

OpenStudy (anonymous):

Should I draw out what I did?

OpenStudy (anonymous):

3/24y-13=x 13+x=3/24y add 13 both sides (13+x)*(24/3)=y multiply by reciprocal

OpenStudy (anonymous):

(13*(24/3)+x*(24/3)=y distribute 24/3

OpenStudy (anonymous):

I kinda see what you're doing... but you lost me

OpenStudy (anonymous):